OpenStudy (anonymous):

Alfred is practicing typing. The first time he tested himself, he could type 22 words per minute. After practicing for a week, he could type 26 words per minute. After two weeks he could type 30 words per minute. Based on this pattern, predict how fast Alfred will be able to type after four weeks of practice.

OpenStudy (anonymous):

x | y ----- 0 22 1 26 2 30 slope (m) = (y2 - y1) / (x2 - x1) (0,22)....x1 = 0 and y1 = 22 (1,26)....x2 = 1 and y2 = 26 slope (m) = (26 - 22) / (1 - 0) slope (m) = 4 y = mx + b (0,22)....x = 0 and y = 22 slope(m) = 4 22 = 4(0) + b 22 = b equation is : y = 4x + 22 f(x) = 4x + 22.....when x(number of weeks) = 4 f(4) = 4(4) + 22 f(4) = 16 + 22 f(4) = 38 4 weeks and he can type 38 words per minute